# Non-Rigid Designators in Modal and Temporal Free Description Logics

### v1

@@ Appendix D Proofs for Section 5
    $\vec{v}=(\vec{w}_{0},\sigma_{0})i_{0}(\vec{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\vec{w}_{m-1},\sigma_{m-1})$ and $i_{m-1}=i$.
    We have $\sigma_{m}=\tau^{\prime}\circ\sigma_{m-1}$ for some $\tau^{\prime}\in\text{Rep}(\vec{w}_{m-1})$. Note that $\rho^{\uparrow}(\vec{u})=\tau^{\prime}\circ\sigma_{m-1}(\rho)(\vec{w}_{m-1}iw_{m})$ and $\rho^{\uparrow}(\vec{v})=\sigma_{m-1}(\rho)(\vec{w}_{m-1})$. Now $\tau^{\prime}\circ\sigma_{m-1}(\rho)$ coincides with $\sigma_{m-1}(\rho)$ on $\vec{w}_{m-1}$. Hence $C\in\tau^{\prime}\circ\sigma_{m-1}(\rho)(\vec{w}_{m-1})$. Hence $\Diamond_{i}C\in\tau^{\prime}\circ\sigma_{m-1}(\rho)(\vec{w}_{m-1}iw_{m})$ since $\tau^{\prime}\circ\sigma_{m-1}$ satisfies condition (R1) for runs in $\mathfrak{Q}^{\prime}$, as required.
@@
    for $\vec{v}^{\prime}=(\vec{w}_{0},\sigma_{0})i_{0}(\vec{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\vec{w}_{m-1},\sigma_{m-1})$ and $i_{m-1}=i$ we have $\vec{v}=\vec{v}^{\prime}i(\vec{w}_{m-1}iw_{m}^{\prime},\sigma_{m}^{\prime})$ for some $(w_{m}^{\prime},\sigma_{m}^{\prime})$ with $\vec{w}_{m-1}iw_{m}^{\prime}\in W^{\prime}$ and $\sigma_{m}^{\prime}=\tau\circ\sigma_{m-1}$ for some $\tau\in\text{Rep}(\vec{w}_{m-1})$. We can then first show that $\Diamond_{i}C\in\rho^{\uparrow}(\vec{v}^{\prime})$ in exactly the same way as in the proof of Point 2. Now we can argue in the same way as in the proof of Point 3 and using (R3) that $\Diamond_{i}C\in\rho^{\uparrow}(\vec{u})$.
For (R2), assume $\Diamond_{i}C\in\rho^{\uparrow}(\vec{u})$. We have to show that there exists $\vec{v}\in W^{\prime\prime}$ such that $\vec{u}R_{i}^{\prime\prime}\vec{v}$ and $C\in\rho^{\uparrow}(\vec{v})$. We have $\rho^{\uparrow}(\vec{u})=\sigma_{m}(\rho)(\vec{w}_{m})$ and distinguish the following three cases:
@@
    $C\in\sigma_{m}(\rho)(\vec{w}_{m})$. Then we are done since then for $\vec{v}=\vec{u}$ we have $C\in\rho^{\uparrow}(\vec{v})$ and we have $\vec{u}R_{i}^{\prime\prime}\vec{v}$.
@@
    for ${\boldsymbol{t}}=\sigma_{m}(\rho)(\vec{w}_{m})$ we have for the proto-run $\rho_{\vec{w}_{m},{\boldsymbol{t}}}\in\mathfrak{R}$ that there is a proto-witness $\vec{w}_{m+1}=v_{\vec{w}_{m},i,C,{\boldsymbol{t}}}\in W^{\prime}$ of the form $\vec{w}_{m}iw_{m+1}$. Then $C\in\rho_{\vec{w}_{m},{\boldsymbol{t}}}(\vec{w}_{m+1})$.
    Then let $\vec{v}=\vec{u}i(\vec{w}_{m+1},\tau\circ\sigma_{m})$ where $\tau$ swaps $\sigma_{m}(\rho)$ with $\rho_{\vec{w}_{m},{\boldsymbol{t}}}$. We have $\vec{u}R_{i}^{\prime\prime}\vec{v}$ and $\rho^{\uparrow}(\vec{v})=\tau\circ\sigma_{m}(\rho)(\vec{w}_{m+1})=\rho_{\vec{w}_{m},{\boldsymbol{t}}}(\vec{w}_{m+1})$. So $C\in\rho^{\uparrow}(\vec{v})$, as required.
@@
    $i_{m-1}=i$ and $\Diamond_{i}C\in\sigma_{m-1}(\rho)(\vec{w}_{m-1})$. If $C\in\sigma_{m-1}(\rho)(\vec{w}_{m-1})$, then $C\in\rho^{\uparrow}(\vec{v})$ for $\vec{v}=(\vec{w}_{0},\sigma_{0})i_{0}(\vec{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\vec{w}_{m-1},\sigma_{m-1})$ and we are done.
    Otherwise we can proceed with $\sigma_{m-1}(\rho)(\vec{w}_{m-1})$ in the same way as we did with $\sigma_{m}(\rho)(\vec{w}_{m})$ in Point 2 and find for $\vec{v}^{\prime}=(\vec{w}_{0},\sigma_{0})i_{0}(\vec{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\vec{w}_{m-1},\sigma_{m-1})$ a $\vec{v}=\vec{v}^{\prime}i(\vec{w}_{m-1}iw_{m}^{\prime},\sigma_{m}^{\prime})$ for some $(w_{m}^{\prime},\sigma_{m}^{\prime})$ with $\vec{w}_{m-1}iw_{m}^{\prime}\in W^{\prime}$ and $\sigma_{m}^{\prime}=\tau\circ\sigma_{m-1}$ for some $\tau\in\text{Rep}(\vec{w}_{m-1})$ such that $C\in\rho^{\uparrow}(\vec{v})$.
@@
Note that there are at most $2^{|\mathsf{con}(C_{0})|}$ types for $C_{0}$. We fix an ordering ${\boldsymbol{t}}_{1},\ldots,{\boldsymbol{t}}_{k}$ of the set of types for $C_{0}$. We use vectors $\vec{x}=(x_{1},\ldots,x_{k})\in(\mathbb{N}\cup\{\infty\})^{k}$ to represent that the type ${\boldsymbol{t}}_{i}$ is satisfied by $x_{i}$-many elements in a world $w$, for $1\leq i\leq n$. Let $|\vec{x}|=\sum_{i=1}^{k}x_{i}$. A quasistate for $C_{0}$ is a vector $\vec{x}=(x_{1},\ldots,x_{k})\in(\mathbb{N}\cup\{\infty\})^{k}$ with $|\vec{x}|>0$ satisfying the following conditions:
@@
    for every $x_{i}>0$ and $\exists u.C\in{\boldsymbol{t}}_{i}$ iff there exists $x_{j}>0$ with such that $C\in{\boldsymbol{t}}_{j}$.
@@
Call a subset $V$ of $W$ $R_{i}$-closed if $v\in V$ whenever $w\in V$ and $wR_{i}v$. Call $V$ $R$-closed if $V$ is $R_{i}$-closed for $1\leq i\leq n$. A run through $(\mathfrak{F},\boldsymbol{q})$ is a partial function $\rho$ mapping worlds $w\in W$ to natural numbers $\rho(w)\leq k$ with $\boldsymbol{q}(w)_{\rho(w)}>0$ such that the domain of $\rho$ is $R$-closed and satisfies the following condition for every $\Diamond_{i}C\in\mathsf{con}(C_{0})$:
@@
  • $\boldsymbol{q}$ is the function from $W$ to the set of quasistates for $C_{0}$ defined by setting $\boldsymbol{q}(w)=(x_{1},\ldots,x_{k})$ with
@@
Observe that $\Delta$ is well-defined and expanding since $|\vec{x}|>0$ for every quasistate $\vec{x}$ and since the domain of each $\rho\in\mathfrak{R}$ is $R$-closed. The following claim is now straightforward from the definition of quasimodels for $C_{0}$.
@@
Now we can easily finish the proof of the proposition by observing that, by (B1), there exists $i\leq k$ with $\boldsymbol{q}(w_{0})_{i}>0$ such that $C_{0}\in{\boldsymbol{t}}_{i}$. Then there exists a $\rho\in\mathfrak{R}$ with $\rho(w_{0})=i$. Hence, by the above claim, $\rho\in C_{0}^{\mathcal{I}_{w_{0}}}$. ∎
@@
Our next aim is to show that there exists a quasimodel for $C_{0}$ of size bounded by a recursive function in $|C_{0}|$ (in particular with $|\boldsymbol{q}(w)|<\infty$ for all $w$). We are going to apply Dickson’s Lemma. For some $k>0$, let $(\mathbb{N}^{k},\leq)$ be the set of $k$-tuples of natural numbers ordered by the natural product ordering: for $\vec{x}=(x_{1},\ldots,x_{k})$ and $\vec{y}=(y_{1},\ldots,y_{k})$ we set $\vec{x}\leq\vec{y}$ if $x_{i}\leq y_{i}$ for $1\leq i\leq k$. A pair $\vec{x},\vec{y}$ with $\vec{x}\leq\vec{y}$ is called an increasing pair. Dickson’s Lemma states every infinite sequence $\vec{x}_{1},\vec{x}_{2},\ldots\in\mathbb{N}^{k}$ contains an increasing pair $\vec{x}_{i_{1}},\vec{x}_{i_{2}}$ with $i_{1}<i_{2}$. In fact, assuming $|\vec{x}_{i}|\leq|\vec{x}_{i+1}|$ for all $i\geq 0$ and given recursive bounds on $|\vec{x}_{1}|$ and $|\vec{x}_{i+1}|-|\vec{x}_{i}|$ one can compute a recursive bound on the length of the longest sequence without any increasing pair (?).
Assume a finite tree-shaped frame $\mathfrak{F}=(W,R_{1},\ldots,R_{n})$ with root $w_{0}$ and worlds of the form (9) is given. Then we say that $\vec{w}$ is the predecessor of $\vec{w}iw\in W$ in $\mathfrak{F}$ and that $\vec{w}iw\in W$ is a successor of $\vec{w}$ in $\mathfrak{F}$. Worlds reachable along a path of predecessors from a world are called ancestors and worlds reachable along a path of successors from a world are called descendants. We set $\vec{w}<\vec{w}^{\prime}$ if $\vec{w}$ is a predecessor of $\vec{w}^{\prime}$ and $W_{\vec{w}}=\{\vec{w}\}\cup\{\vec{v}\in W\mid\vec{w}<\vec{v}\}$.
We now show that one can reduce quasimodels for $C_{0}$ to quasimodels for $C_{0}$ not having increasing pairs $|\boldsymbol{q}(\vec{w})|,|\boldsymbol{q}(\vec{w}^{\prime})|$ with $\vec{w}<\vec{w}^{\prime}$ and satisfying recursive bounds for $|\boldsymbol{q}(w_{0})|$, $|\boldsymbol{q}(\vec{w}iw)|-|\boldsymbol{q}(\vec{w})|$, and the outdegree of each $\vec{w}$. We then obtain a recursive bound on the size of such quasimodels from Dickson’s Lemma.
@@
    if $v$ is a successor of $w$, then $|\boldsymbol{q}(v)|\leq|\boldsymbol{q}(w)|+2^{|C_{0}|}$;
@@
    if $w<v$, then $\boldsymbol{q}(w)\not\leq\boldsymbol{q}(v)$;
@@
    the outdgree of $w$ in $\mathfrak{F}$ is bounded by $|\boldsymbol{q}(w)|\times|C_{0}|$.
@@
Assume that a quasimodel $\mathfrak{Q}$ for $C_{0}$ with $\mathfrak{Q}=(\mathfrak{F},\boldsymbol{q},\mathfrak{R})$ and $\mathfrak{F}=(W,R_{1},\ldots,R_{n})$ with root $w_{0}$ is given. We may assume that every run $\rho\in\mathfrak{R}$ has a root $w\in W$ in the sense that $W_{w}=\text{dom}(\rho)$. We let $\mathfrak{R}_{w}$ denote the set of runs with root $w$. We introduce a few rules that allow us to modify $\mathfrak{Q}$:
Minimize root. Assume $\boldsymbol{q}(w_{0})=(x_{1},\ldots,x_{k})$. Pick for any $x_{i}>0$ a single $\rho_{i}\in\mathfrak{R}$ with $\rho_{i}(w_{0})=i$. Now construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining
@@
  • $\boldsymbol{q}^{\prime}(w_{0})=(\min\{x_{1},1\},\ldots,\min\{x_{k},1\})$ and $\boldsymbol{q}^{\prime}(w)=\boldsymbol{q}(w)$ for all $w\not=w_{0}$,
  • $\mathfrak{R}^{\prime}$ as the set of selected runs $\rho_{i}$, all runs in $\mathfrak{R}\setminus\mathfrak{R}_{w_{0}}$, and the restrictions of all remaining $\rho\in\mathfrak{R}_{w_{0}}$ to $W_{w}$ for $w$ a successor of $w_{0}$.
@@
Minimize non-root: Assume $v$ is a successor of $w$ and $\boldsymbol{q}(v)=(x_{1},\ldots,x_{k})$. Call $i\leq k$ increasing at $v$ if there exists $\rho_{i}\in\mathfrak{R}_{v}$ such that $\rho_{i}(v)=i$. Select for any such $i$ such a $\rho_{i}\in\mathfrak{R}_{v}$. We construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining
@@
  • $\boldsymbol{q}^{\prime}(v)=(x_{i}^{\prime},\ldots,x_{k}^{\prime})$ with
    $x_{i}^{\prime}=|\{\rho\in\mathfrak{R}\mid w\in\text{dom}(\rho),\rho(v)=i\}|+1$
    if $i$ is increasing at $v$ and $x_{i}^{\prime}=x_{i}$ otherwise. Set $\boldsymbol{q}^{\prime}(w^{\prime})=\boldsymbol{q}(w^{\prime})$ for all $w^{\prime}\not=v$.
  • $\mathfrak{R}^{\prime}$ as the set of selected runs $\rho_{i}$, all runs in $\mathfrak{R}\setminus\mathfrak{R}_{v}$, and the restrictions of all remaining $\rho\in\mathfrak{R}_{v}$ to $W_{v^{\prime}}$ for $v^{\prime}$ a successor of $v$.
It is easy to see that $\mathfrak{Q}^{\prime}$ is a quasimodel for $C_{0}$ satisfying Condition 2 for $w$ and $v$.
Drop interval 1: Consider $\vec{w}=\vec{w}^{\prime}iw\in W$ and $\vec{v}=\vec{v^{\prime}}jv\in W$ with $\vec{w}<\vec{v}$ and
  • $\boldsymbol{q}(\vec{w})=(x_{1},\ldots,x_{k})$;
  • $\boldsymbol{q}(\vec{v})=(y_{1},\ldots,y_{k})$;
  • $x_{i}\leq y_{i}$ for all $i\leq k$.
Now construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining

### v2

@@ D.1 Proof of Theorem 11
    $\boldsymbol{v}=(\boldsymbol{w}_{0},\sigma_{0})i_{0}(\boldsymbol{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\boldsymbol{w}_{m-1},\sigma_{m-1})$ and $i_{m-1}=i$.
    We have $\sigma_{m}=\tau^{\prime}\circ\sigma_{m-1}$ for some $\tau^{\prime}\in\textrm{Rep}(\boldsymbol{w}_{m-1})$. Note that $\rho^{\uparrow}(\boldsymbol{u})=\tau^{\prime}\circ\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1}iw_{m})$ and $\rho^{\uparrow}(\boldsymbol{v})=\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1})$. Now $\tau^{\prime}\circ\sigma_{m-1}(\rho)$ coincides with $\sigma_{m-1}(\rho)$ on $\boldsymbol{w}_{m-1}$. Hence, $C\in\tau^{\prime}\circ\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1})$. Thus, as required, $\Diamond_{i}C\in\tau^{\prime}\circ\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1}iw_{m})$ since $\tau^{\prime}\circ\sigma_{m-1}$ satisfies condition (R1) for runs in $\mathfrak{Q}^{\prime}$.
@@
    for $\boldsymbol{v}^{\prime}=(\boldsymbol{w}_{0},\sigma_{0})i_{0}(\boldsymbol{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\boldsymbol{w}_{m-1},\sigma_{m-1})$ and $i_{m-1}=i$ we have $\boldsymbol{v}=\boldsymbol{v}^{\prime}i(\boldsymbol{w}_{m-1}iw_{m}^{\prime},\sigma_{m}^{\prime})$ for some $(w_{m}^{\prime},\sigma_{m}^{\prime})$ with $\boldsymbol{w}_{m-1}iw_{m}^{\prime}\in W^{\prime}$ and $\sigma_{m}^{\prime}=\tau\circ\sigma_{m-1}$ for some $\tau\in\textrm{Rep}(\boldsymbol{w}_{m-1})$.
    We can then first show that $\Diamond_{i}C\in\rho^{\uparrow}(\boldsymbol{v}^{\prime})$ in exactly the same way as in the proof of Point 2. Now we can argue in the same way as in the proof of Point 3 and using (R3) that $\Diamond_{i}C\in\rho^{\uparrow}(\boldsymbol{u})$.
For (R2), assume $\Diamond_{i}C\in\rho^{\uparrow}(\boldsymbol{u})$. We have to show that there exists $\boldsymbol{v}\in W^{\prime\prime}$ such that $\boldsymbol{u}R_{i}^{\prime\prime}\boldsymbol{v}$ and $C\in\rho^{\uparrow}(\boldsymbol{v})$. We have $\rho^{\uparrow}(\boldsymbol{u})=\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ and distinguish the following three cases.
@@
    $C\in\sigma_{m}(\rho)(\boldsymbol{w}_{m})$. Then we are done since, for $\boldsymbol{v}=\boldsymbol{u}$, we have $C\in\rho^{\uparrow}(\boldsymbol{v})$ and $\boldsymbol{u}R_{i}^{\prime\prime}\boldsymbol{v}$.
@@
    For ${\boldsymbol{t}}=\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ we have for the proto-run $\rho_{\boldsymbol{w}_{m},{\boldsymbol{t}}}\in\mathfrak{R}$ that there is a proto-witness $\boldsymbol{w}_{m+1}=v_{\boldsymbol{w}_{m},i,C,{\boldsymbol{t}}}\in W^{\prime}$ of the form $\boldsymbol{w}_{m}iw_{m+1}$. Then $C\in\rho_{\boldsymbol{w}_{m},{\boldsymbol{t}}}(\boldsymbol{w}_{m+1})$.
    Then let $\boldsymbol{v}=\boldsymbol{u}i(\boldsymbol{w}_{m+1},\tau\circ\sigma_{m})$, where $\tau$ swaps $\sigma_{m}(\rho)$ with $\rho_{\boldsymbol{w}_{m},{\boldsymbol{t}}}$. We have $\boldsymbol{u}R_{i}^{\prime\prime}\boldsymbol{v}$ and $\rho^{\uparrow}(\boldsymbol{v})=\tau\circ\sigma_{m}(\rho)(\boldsymbol{w}_{m+1})=\rho_{\boldsymbol{w}_{m},{\boldsymbol{t}}}(\boldsymbol{w}_{m+1})$. So, $C\in\rho^{\uparrow}(\boldsymbol{v})$, as required.
@@
    $i_{m-1}=i$ and $\Diamond_{i}C\in\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1})$.
    If $C\in\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1})$, then $C\in\rho^{\uparrow}(\boldsymbol{v})$ for $\boldsymbol{v}=(\boldsymbol{w}_{0},\sigma_{0})i_{0}(\boldsymbol{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\boldsymbol{w}_{m-1},\sigma_{m-1})$, and we are done.
    Otherwise, we can proceed with $\sigma_{m-1}(\rho)(\boldsymbol{w}_{m-1})$ in the same way as we did with $\sigma_{m}(\rho)(\boldsymbol{w}_{m})$ in Point 2 and find for $\boldsymbol{v}^{\prime}=(\boldsymbol{w}_{0},\sigma_{0})i_{0}(\boldsymbol{w}_{1},\sigma_{1})i_{1}\cdots i_{m-2}(\boldsymbol{w}_{m-1},\sigma_{m-1})$ a $\boldsymbol{v}=\boldsymbol{v}^{\prime}i(\boldsymbol{w}_{m-1}iw_{m}^{\prime},\sigma_{m}^{\prime})$ for some $(w_{m}^{\prime},\sigma_{m}^{\prime})$ with $\boldsymbol{w}_{m-1}iw_{m}^{\prime}\in W^{\prime}$ and $\sigma_{m}^{\prime}=\tau\circ\sigma_{m-1}$ for some $\tau\in\textrm{Rep}(\boldsymbol{w}_{m-1})$ such that $C\in\rho^{\uparrow}(\boldsymbol{v})$.
@@ D.2 Proof of Theorem 12
Note that there are at most $2^{|\mathsf{con}(C_{0})|}$ types for $C_{0}$. We fix an ordering ${\boldsymbol{t}}_{1},\ldots,{\boldsymbol{t}}_{k}$ of the set of types for $C_{0}$. We use vectors $\boldsymbol{x}=(x_{1},\ldots,x_{k})\in\mathbb{N}_{\infty}^{k}$ to represent that the type ${\boldsymbol{t}}_{i}$ is satisfied by $x_{i}$-many elements in a world $w$, for $1\leq i\leq n$. Let $|\boldsymbol{x}|=\sum_{i=1}^{k}x_{i}$. A quasistate for $C_{0}$ is a vector $\boldsymbol{x}=(x_{1},\ldots,x_{k})\in\mathbb{N}_{\infty}^{k}$ with $|\boldsymbol{x}|>0$ satisfying the following conditions:
@@
    for every $x_{i}>0$ and $\exists u.C\in{\boldsymbol{t}}_{i}$ iff there exists $x_{j}>0$ such that $C\in{\boldsymbol{t}}_{j}$.
@@
Call a subset $V$ of $W$ $R_{i}$-closed if $v\in V$ whenever $w\in V$ and $wR_{i}v$. Call $V$ $R$-closed if $V$ is $R_{i}$-closed for $1\leq i\leq n$. A run through $(\mathfrak{F},\boldsymbol{q})$ is a partial function $\rho$ mapping worlds $w\in W$ to natural numbers $\{1,\dots,k\}$ with $\boldsymbol{q}(w)_{\rho(w)}>0$ such that the domain of $\rho$ is $R$-closed and satisfies the following condition for every $\Diamond_{i}C\in\mathsf{con}(C_{0})$:
@@
  • $\boldsymbol{q}$ is the function from $W$ to the set of quasistates for $C_{0}$ defined by setting $\boldsymbol{q}(w)=(x_{1},\ldots,x_{k})\in\mathbb{N}_{\infty}^{k}$ with
@@
Observe that $\Delta$ is well-defined and expanding since $|\boldsymbol{x}|>0$ for every quasistate $\boldsymbol{x}$ and since the domain of each $\rho\in\mathfrak{R}$ is $R$-closed. The following claim is now straightforward from the definition of quasimodels for $C_{0}$.
@@
Now we can easily finish the proof of Lemma 27 by observing that, by (B1), there exists $i\leq k$ with $\boldsymbol{q}(w_{0})_{i}>0$ such that $C_{0}\in{\boldsymbol{t}}_{i}$. Then there exists a $\rho\in\mathfrak{R}$ with $\rho(w_{0})=i$. Hence, by Claim 6, $\rho\in C_{0}^{\mathcal{I}_{w_{0}}}$. ∎
@@
Our next aim is to show that there exists a quasimodel for $C_{0}$ of size bounded by a recursive function in $|C_{0}|$ (in particular with $|\boldsymbol{q}(w)|<\infty$ for all $w$). We are going to apply Dickson’s Lemma. For some $k>0$, let $(\mathbb{N}^{k},\leq)$ be the set of $k$-tuples of natural numbers ordered by the natural product ordering: for $\boldsymbol{x}=(x_{1},\ldots,x_{k})$ and $\boldsymbol{y}=(y_{1},\ldots,y_{k})$ we set $\boldsymbol{x}\leq\boldsymbol{y}$ if $x_{i}\leq y_{i}$ for $1\leq i\leq k$. A pair $\boldsymbol{x},\boldsymbol{y}$ with $\boldsymbol{x}\leq\boldsymbol{y}$ is called an increasing pair. Dickson’s Lemma states every infinite sequence $\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots\in\mathbb{N}^{k}$ contains an increasing pair $\boldsymbol{x}_{i_{1}},\boldsymbol{x}_{i_{2}}$ with $i_{1}<i_{2}$. In fact, assuming $|\boldsymbol{x}_{i}|\leq|\boldsymbol{x}_{i+1}|$ for all $i\geq 0$ and given recursive bounds on $|\boldsymbol{x}_{1}|$ and $|\boldsymbol{x}_{i+1}|-|\boldsymbol{x}_{i}|$ one can compute a recursive bound on the length of the longest sequence without any increasing pair (?).
Assume a finite tree-shaped frame $\mathfrak{F}=(W,R_{1},\ldots,R_{n})$ with root $w_{0}$ and worlds of the form (9) is given. Then we say that $\boldsymbol{w}$ is the predecessor of $\boldsymbol{w}iw\in W$ in $\mathfrak{F}$ and that $\boldsymbol{w}iw\in W$ is a successor of $\boldsymbol{w}$ in $\mathfrak{F}$. Worlds reachable along a path of predecessors from a world are called ancestors and worlds reachable along a path of successors from a world are called descendants. We set $\boldsymbol{w}<\boldsymbol{v}$ if $\boldsymbol{w}$ is a predecessor of $\boldsymbol{v}$ and denote $W_{\boldsymbol{w}}=\{\boldsymbol{w}\}\cup\{\boldsymbol{v}\in W\mid\boldsymbol{w}<\boldsymbol{v}\}$.
We now show that one can reduce quasimodels for $C_{0}$ to quasimodels for $C_{0}$ not having increasing pairs $\boldsymbol{q}(\boldsymbol{w}),\boldsymbol{q}(\boldsymbol{w}^{\prime})$ with $\boldsymbol{w}<\boldsymbol{w}^{\prime}$ and satisfying recursive bounds for $|\boldsymbol{q}(w_{0})|$, $|\boldsymbol{q}(\boldsymbol{w}iw)|-|\boldsymbol{q}(\boldsymbol{w})|$ and the outdegree of each $\boldsymbol{w}$. We then obtain a recursive bound on the size of such quasimodels from Dickson’s Lemma.
@@
    if $\boldsymbol{v}$ is a successor of $\boldsymbol{w}$, then $|\boldsymbol{q}(\boldsymbol{v})|\leq|\boldsymbol{q}(\boldsymbol{w})|+2^{|C_{0}|}$;
@@
    if $\boldsymbol{w}<\boldsymbol{v}$, then $\boldsymbol{q}(\boldsymbol{w})\not\leq\boldsymbol{q}(\boldsymbol{v})$;
@@
    the outdgree of $\boldsymbol{w}$ in $\mathfrak{F}$ is bounded by $|\boldsymbol{q}(\boldsymbol{w})|\times|C_{0}|$.
@@
Assume that a quasimodel $\mathfrak{Q}$ for $C_{0}$ with $\mathfrak{Q}=(\mathfrak{F},\boldsymbol{q},\mathfrak{R})$ and $\mathfrak{F}=(W,R_{1},\ldots,R_{n})$ with root $w_{0}$ is given. We may assume that every run $\rho\in\mathfrak{R}$ has a root $\boldsymbol{w}\in W$ in the sense that $W_{\boldsymbol{w}}=\text{dom}(\rho)$. We let $\mathfrak{R}_{\boldsymbol{w}}$ denote the set of runs with root $\boldsymbol{w}$. We introduce a few rules that allow us to modify $\mathfrak{Q}$.
Minimize root: Assume $\boldsymbol{q}(w_{0})=(x_{1},\ldots,x_{k})$. Pick, for any $x_{i}>0$, a single $\rho_{i}\in\mathfrak{R}$ with $\rho_{i}(w_{0})=i$. Now construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining
@@
  • $\boldsymbol{q}^{\prime}(w_{0})=(\min\{x_{1},1\},\ldots,\min\{x_{k},1\})$ and $\boldsymbol{q}^{\prime}(\boldsymbol{w})=\boldsymbol{q}(\boldsymbol{w})$ for all $\boldsymbol{w}\neq w_{0}$,
  • $\mathfrak{R}^{\prime}$ as the set of selected runs $\rho_{i}$, all runs in $\mathfrak{R}\setminus\mathfrak{R}_{w_{0}}$, and the restrictions of all remaining $\rho\in\mathfrak{R}_{w_{0}}$ to $W_{\boldsymbol{w}}$ for $\boldsymbol{w}$ a successor of $w_{0}$.
@@
Minimize non-root: Assume $\boldsymbol{v}$ is a successor of $\boldsymbol{w}$ and $\boldsymbol{q}(\boldsymbol{v})=(x_{1},\ldots,x_{k})$. Call $i\leq k$ increasing at $\boldsymbol{v}$ if there exists $\rho_{i}\in\mathfrak{R}_{\boldsymbol{v}}$ such that $\rho_{i}(\boldsymbol{v})=i$. Select, for any $i$ increasing at $\boldsymbol{v}$, such a $\rho_{i}\in\mathfrak{R}_{\boldsymbol{v}}$. We construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining
@@
  • $\boldsymbol{q}^{\prime}(\boldsymbol{v})=(x_{i}^{\prime},\ldots,x_{k}^{\prime})$ with
    $x_{i}^{\prime}=|\{\rho\in\mathfrak{R}\mid\boldsymbol{w}\in\text{dom}(\rho),\rho(\boldsymbol{v})=i\}|+1$
    if $i$ is increasing at $\boldsymbol{v}$ and $x_{i}^{\prime}=x_{i}$ otherwise. Set $\boldsymbol{q}^{\prime}(\boldsymbol{w}^{\prime})=\boldsymbol{q}(\boldsymbol{w}^{\prime})$ for all $\boldsymbol{w}^{\prime}\neq\boldsymbol{v}$.
  • $\mathfrak{R}^{\prime}$ as the set of selected runs $\rho_{i}$, all runs in $\mathfrak{R}\setminus\mathfrak{R}_{\boldsymbol{v}}$, and the restrictions of all remaining $\rho\in\mathfrak{R}_{\boldsymbol{v}}$ to $W_{\boldsymbol{v}^{\prime}}$ for $\boldsymbol{v}^{\prime}$ a successor of $\boldsymbol{v}$.
It is easy to see that $\mathfrak{Q}^{\prime}$ is a quasimodel for $C_{0}$ satisfying Condition 2 for $\boldsymbol{w}$ and $\boldsymbol{v}$.
Drop interval 1: Consider $\boldsymbol{w}=\boldsymbol{w}^{\prime}iw\in W$ and $\boldsymbol{v}=\boldsymbol{v^{\prime}}jv\in W$ with $\boldsymbol{w}<\boldsymbol{v}$ and $\boldsymbol{q}(\boldsymbol{w})\leq\boldsymbol{q}(\boldsymbol{v})$. Now construct an updated $\mathfrak{Q}^{\prime}=(\mathfrak{F}^{\prime},\boldsymbol{q}^{\prime},\mathfrak{R}^{\prime})$ by defining